Two charges are located on the positive X-axis of a

coordinate system. If and are 5cm

and 7cm from the origin (X=0). What is the force

exerted by & on a third charge ; if

A. is placed at 3cm from the origin?




1
Expert's answer
2021-03-18T13:14:42-0400

Consider the diagram below. We can assume that Q1 = 6nC, Q2 = -9nC and Q3 = 5nC since the question is incomplete


"F_{31}=\frac{kq_1q_3}{r^2}=\frac{9 \times 10^9 \times(6 \times10^{-9})(5 \times10^{-9})}{(2 \times10^{-4})^2}=6.75 N"


"F_{32}=\frac{kq_2q_3}{r^2}=\frac{9 \times 10^9 \times(-9 \times10^{-9})(5 \times10^{-9})}{(4 \times10^{-4})^2}=-2.53125 N"


Net force on Q3 ="F_{32}+F_{31}=6.75+-2.53125=4.21875 N"